3 Check Internal Q-scan & T-scan Interlock system works fine –N2 air flow –Temperature HV –HV modules works –Noise shape seems fine for all PMTs. Laser data –All raw PMTs shape for laser signal are fine. –We took laser data by DAQ. Laser signal Adc tdc0tdc1 Q-scan T-scan Laser signal
5 Time in Change MTM timing by v124(ch7) –Fine Delay 132 → 116 (8ns) Time in at 12/28!! ch1: BBC GTM ch2: Discriminator output ~60ns ~68ns Change delay But, BBLL1 trigger has problem… Above figure is obtained by ZDCLL1 trigger
6 Online monitor Trigger is ZDCLL1 BBLL1 has problem. We can NOT take data with BBLL1 trigger –prescale of BBLL1 set to 999999 –LL1 expert is working Now, we are taking data with ZDCLL1 trigger. The missing plots on BBC monitor is due to BBCLL1. BBC is working fine!!
7 ADC distribution Some PMT channels have overflow –Same HV with Run9 –Fraction of overflow is less 0.1% (see attached file “ hvopt_299055.pdf ”) ch0 ch8
8 PmtGain (pC/MIP) Calibration have done (Run298917) –Same HV with RUN9 –Note: ZDCLL1 trigger events was used PmtGain@Run10 is typically 15% larger than @end of Run9 –Large fluctuation Need new optimized HV for RUN10 @Run10 @end of Run9 (Run10–Run9)/Run9 x100 ch0
9 Determination of HV Gain Curve was Obtained –For 6 settings -100, -50, 0, +25, +50, +75V Optimized HV was obtained –Inner ring :30pC/MIP HV –mid&outter: 40pC/MIP HV ADC distribution Gain curve Determined Optimized HV pC / MIP Used pC/MIP@Run10 as normalization point for blue curve
10 Operational HV RUN10 for +- field (30, 40pC/MIP HV) HV_BB_S-1 -1649.7 HV_BB_S-2 -1914.2 HV_BB_S-3 -1676.3 HV_BB_S-4 -2134.0 HV_BB_S-5 -2029.2 HV_BB_S-6 -1485.4 HV_BB_S-7 -1631.7 HV_BB_S-8 -1957.3 HV_BB_N-1 -1603.1 HV_BB_N-2 -2400.0 HV_BB_N-3 -1716.2 HV_BB_N-4 -1868.6 HV_BB_N-5 -1529.9 HV_BB_N-6 -2250.4 HV_BB_N-7 -1901.1 HV_BB_N-8 -1700.1 The operational HV was calculated by using pC/MIP obtained from #298917 S3,S7,N3,N7(inner) are applied 30pC/MIP HV, others are applied 40pC/MIP HV. Valid 2009/12/31 06:45~, #299319~
11 Time table New HV applied 2009/12/31, 06:45~ #RUN299319~ LUT update 2010/01/01,20:36~ #RUN299480~ DB update 2010/01/01 Valid:2009/12/31, 06:45~ MTM timing turning 116 112 2010/01/02, 20:07~ #RUN299603~ MTM timing turning 132 116 2009/12/28, 09:59~ RUN298902~ BBLL1 fixed 2009/12/31 #RUN299317~ First Collision 2009/12/27 TIME Happy new year !!
12 Outputs #RUN299607 working fine!! ch0 Q-scanT-scan
13 FEM board problem 2010/01/08 –Complete feed, LV recycle could not solve 2010/01/09 –IR access –exchange fuse on FEM board Fixed this problem
14 First PHYSICS Run300476 –Start: 2010/01/10, 01:36:52~ –Fixed FEM problem –Everything FINE!
17 ADC distribution ch26 ch121 Channels which have relatively larger fraction of overflow Full range
18 Expected edge of ADC @new HV Maximum #MIP is obtained from Run298917 –Weighted mean of top 0.5% I guess, weighted mean does not work well for the ch which have overflow. ch8 red: pedestal Blue: weighted mean of top 0.5%
19 To do Need comments and suggestion!! –IF you need bbctree, location of tree is blow /direct/phenix+subsys+bbc/phnxbb/Y10DB/tmp/BbcCalib_29 8917 @rcf (sorry for inconvenience…) * still have not applied new HV. to take data with new HV –Check MIP & overflow After solved BBLL1 problem, I will do calibration again by using BBLL1 triggered event to obtain LUT and DB constants. （ I might also do timing tuning of MTM ） 12/ 29
20 Status New HV is applied –MIP & overflow: within acceptable region LUT & DB is updated –Valid range DB: 2009/12/31, 06:45~ (#RUN299317~) LUT: 2010/01/01, 20:36~ (#RUN299480~) BBLL1 problem fixed Fine turning of MTM timing –116 112 (Fine Delay) 2010/01/04